2. The following genotypes were observed in a population of grizzly bears in Yellowstone National Park. AA1 13 bears The genotypic frequencies expected in this population are AA1 0.436 Use the x2 test to determine if the genotypic frequencies observed match those expected. Be sure to show your work when calculating x2 and clearly present: AA2 7 bears AZA2 12 bears AA2 0.449 A?A2 0.115 a) Your calculated x value b) the number of degrees of freedom for this problem c) the critical x2 value from the table you to which you are comparing your calculated x value sion, either the observed values are consistent with the expected values or they d) your conclu are not.
Answer #1

Answer:

2a).

Observed genotypic frequencies:

A1A1 = 13/32=0.406

A1A2 = 7/32=0.219

A2A2 = 12/32= 0.375

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

A1A1

0.406

0.436

-0.03

0.001

0.002

A1A2

0.219

0.449

-0.23

0.053

0.118

A2A2

0.375

0.115

0.26

0.068

0.588

1

1

0.708

X^2 value = 0.708

b).

Degrees of freedom = number of genotype – 1

Df = 3-1=2

c).

Ther critical X^2 value is 5.99 for df of 2.

d).

The observed values are consistent with the expected values. Because, the X^value of 0.708 is less than the critical value of 5.99. So we can accept the null hypothesis.